OpenStudy (nairatrebilus):

Is the function f(x) = x^(–4) + cos(–2x) even, odd, or neither?

OpenStudy (nairatrebilus):

@phi @mathstudent55

OpenStudy (kirbykirby):

If \(f(x) \) is even, then \(f(-x)=f(x)\) If \(f(x) \) is odd, then \(f(-x)=-f(x)\).

OpenStudy (kirbykirby):

So try see what you get from \(f(-x)=(-x)^{-4}+\cos(-2(-x))\)

OpenStudy (nairatrebilus):

I just realized that cos(-x)=cos(x). Thanks

OpenStudy (kirbykirby):

:)